OpenStudy (anonymous):

What exponential function represents the data in the table? x f(x) 2 25 3 125 4 625 (1 point) f(x) = x4 + 9 f(x) = 4x + 9 f(x) = x5 f(x) = 5x

OpenStudy (campbell_st):

substitute the x value and see what happens choice 5 f(x) = x^5 so if x = 2 (from the table) f(2) = 2^5 = 32 doesn't match the table value of 25.... so eliminate choice 3 try choice 4 f(2) = 5(2) = 10 doesn't match the table so eliminate choice 4 leaves you 1 and 2... so when you substitute x = 2 which gives an answer of 25...